OpenStudy (anonymous):

how to do logarithism (?) problem 90=120e^-0.01/t

OpenStudy (tkhunny):

Divide by 120 Introduce the logarithm.

OpenStudy (anonymous):

so log 3/4 = e^-.01

OpenStudy (anonymous):

oops that e^-.01/t

OpenStudy (tkhunny):

You almost had it. \(3/4 = e^{-0.01/t}\) Introduce the logarithm \(ln(3/4) = -0.01/t\) It's almost done.

OpenStudy (anonymous):

I multiply then right? the ln(3/4) by -.01

OpenStudy (anonymous):

If I multiply I get .0028768207 which is not the right answer

OpenStudy (anonymous):

nevermind I'm a dork I have to multiply t by both sides since it was in the denominator then dividee by ln(3/4)....duh.. Thanks for your help I need to become OCD LOL

OpenStudy (tkhunny):

Well, that was an interesting conclusion. :-)